# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ Abstract
Why does a phenomenon occur? Addressing this question is central to most scientific inquiries based on empirical observations, and often heavily relies on simulations of scientific models. As models become more intricate, deciphering the causes behind these phenomena in high-dimensional spaces of interconnected variables becomes increasingly challenging. Causal machine learning may assist scientists in the discovery of relevant and interpretable patterns of causation in simulations. We introduce Targeted Causal Reduction (TCR), a method for turning complex models into a concise set of causal factors that explain a specific target phenomenon. We derive an information theoretic objective to learn TCR from interventional data or simulations and propose algorithms to optimize this objective efficiently. TCR’s ability to generate interpretable high-level explanations from complex models is demonstrated on toy and mechanical systems, illustrating its potential to assist scientists in the study of complex phenomena in a broad range of disciplines.
@@ 1 INTRODUCTION
Mathematical models are crucial in science for representing real-world systems and serve various purposes, including generating etiological explanations—identifying the causes of specific phenomena. For instance, general circulation models address the causes of global warming (Grassl, 2000), while computational models of the brain investigate the origins of pathologies (Breakspear, 2017; Deco and Kringelbach, 2014). These examples illustrate the increasing complexity of scientific models, designed to faithfully capture the large number of mechanisms at play in these systems. However, this complexity comes at a cost: expanding parameter spaces and heightened computational demands. This trend, in turn, impacts the ability to generate high-level explanations, understandable by scientists and decision makers (Reichstein et al., 2019; Safavi et al., 2023). This is analogous to challenges of explainable AI: explaining a system can be difficult even when its mechanisms are perfectly known.
Figure 1: Targeted Causal Reduction. (a) Example targeted model reduction: a model of the dynamics of a system of point masses can be reduced to the trajectory of its center of gravity. (b) Principle of TCR.
Consider the system of point masses depicted in Fig. 0(a). Its trajectory can be accurately predicted by simulating the coupled equations of motion of individual point masses. Suppose, we are interested in one property of this system’s trajectory that takes the form of one single scalar variable $Y$, that we call the target variable. Let us take $Y$ to be the horizontal speed of the system’s center of gravity. We would like to discover its “causes”, meaning which properties of the system influence $Y$ and the mechanisms mediating this influence. In principle, it is possible to derive these from this “microscopic” model. However, this set is large as it includes all internal and external forces exerted on individual points. In contrast, the horizontal motion of the center of gravity can be described by a single scalar differential equation, subject to a single net force—the sum of all horizontal components of external forces.
This illustrates the goal we want to achieve, depicted in Fig. 0(b). Starting from a complex low-level model described by a vector of variables $\mathbf{X}$ and a target scalar variable $Y$ derived from it (such that $Y=\tau_{0}(\mathbf{X})$, for some function $\tau_{0}$), we want to learn a transformation that maps $\mathbf{X}$ to fewer variables forming a high-level model. This high-level model should capture the causal influences on $Y$ in a concise and interpretable way, while remaining consistent with the low-level model. This notion of consistency is defined using structural causal models (SCM) (Pearl, 2009) and interventions on them. The variables of a low-level SCM are mapped to variables of a high-level SCM through a transformation $\tau$, while low-level interventions are correspondingly mapped to high-level interventions through another transformation $\omega$.
Ideally, transformations should satisfy the equivalence of the following two paths: (1) intervening on the low-level model and mapping the outcome to the high-level with $\tau$, or (2) first mapping the low-level model to the high-level model with $\tau$ and then applying the high-level intervention derived from the low-level one using $\omega$. A transformation between models satisfying such a consistency constraint is called exact (Rubenstein et al., 2017; Beckers and Halpern, 2019). However, in practice, one may only be able to achieve it approximately (Beckers et al., 2020).
In this work, we use the Kullback-Leibler divergence to formulate the above interventional consistency constraint as a learning objective. Starting from a set of interventional distributions of a given low-level SCM with target variable $Y$, we learn the transformation and the high-level SCM that form a causal explanation for $Y$. We propose the Targeted Causal Reduction algorithm to minimize the consistency objective by stochastic gradient descent and apply it to synthetic and scientific models to demonstrate our approach.
@@ 2.1 Structural causal models
Causal dependencies between multiple variables can be described using Structural Causal Models (SCM) (Peters et al., 2017).
@@
  • a directed acyclic graph $\mathcal{G}$ with $n$ vertices,
@@
  • a joint distribution $P_{\mathbf{U}}$ over the exogenous variables $\{U_{j}\}_{j\leq n}$, assumed jointly independent.
For each value of the exogenous variables, $\mathbb{S}$ leads to a unique solution for the vector of so-called endogenous variables $\mathbf{X}=[X_{1},\dots,X_{n}]^{\top}$, such that the distribution $P_{\mathbf{U}}$ entails a well-defined joint distribution over the endogenous variables $P(\mathbf{X})$. Interventions in SCMs involve replacing one or more structural equations, potentially modifying exogenous distributions, and adding or removing arrows in the original graph to reflect changes in dependencies between variables. An intervention transforms the original model $\mathcal{M}=(\mathcal{G},\mathbb{S},P_{\mathbf{U}})$ into an intervened model $\mathcal{M}^{(\mathbf{i})}=(\mathcal{G}^{(\mathbf{i})},\mathbb{S}^{(\mathbf{i})},P_{\mathbf{U}}^{(\mathbf{i})})$, where $\mathbf{i}$ is the vector parameterizing the intervention. The base probability distribution of the unintervened model is denoted $P_{\mathcal{M}}^{(0)}(\mathbf{X})$ or simply $P_{\mathcal{M}}(\mathbf{X})$ and the interventional distribution associated with $\mathcal{M}^{(\mathbf{i})}$ is denoted $P_{\mathcal{M}}^{(\mathbf{i})}(\mathbf{X})$. When there is no ambiguity, we drop the subscript of the base model $\mathcal{M}$.
While standard $do$-interventions set a structural equation to a constant, they may not correspond to plausible changes of the real-world system being modeled. Instead, soft interventions modify an equation while keeping the set of parents unchanged. Large classes of soft interventions can be designed to match domain knowledge (Besserve and Schölkopf, 2022). Notably, the class of shift interventions modify the structural equation of endogenous variable $l$ through shifting it by a scalar parameter $i$
@@
2.2 Scientific models and simulations
We use the term scientific model to refer to a generative model that relies on a set of equations to represent a phenomenon. What distinguishes such models from generative models in machine learning is their decomposability into elementary functions within parametric classes, encoding domain-specific knowledge about the mechanisms being investigated. We argue that simulators of scientific models can often be expressed as SCMs. This notably includes Ordinary (ODE) and Stochastic Differential Equations (SDE), whose simulation on a discretized grid is performed through explicit numerical schemes. For example, let $F$ be a smooth scalar function and $W$ the standard Brownian motion, the SDE with initial value $X(0)=x_{0}$ and dynamics
$dX=F(X(t))dt+\sigma_{W}\cdot dW\,,$
can be simulated over a grid of fixed time step $\Delta t$ using the Euler–Maruyama method (Sauer, 2013), leading to approximate values $X_{k}\approx X(k\Delta t)$ following the scheme
$\begin{cases}X_{1}\coloneqq x_{0}+\Delta t\cdot F(x_{0})+U_{1}&\\ X_{k+1}\coloneqq X_{k}+\Delta t\cdot F(X_{k})+U_{k+1}\,,&k>1\,,\end{cases}$
where $\{U_{k}\}$ are independent Gaussian variables (see App. A.1 for more details). These sets of equations, spanning all time points, can thus be viewed as a valid SCM with a (linear) DAG structure. Generalization of this view to multidimensional ODE/SDEs as well as Partial Differential Equations is possible (see App. A.1 for more details). Soft interventions on such models, such as shift interventions, can be conceived as modeling perturbations of the real-world system. See App. A.2 for an example system of point masses.
A simulator can thus be seen as a low-level causal model, from which samples of unintervened and intervened distributions can be queried. Those samples will be used to learn a reduction to a high-level causal model.
@@ 2.3 Causal Model Reductions (CMR)
We consider as CMR any (possibly approximate) mapping from a low-level SCM $\mathcal{L}$ to a simpler high-level SCM $\mathcal{H}$. An example is Causal Feature Learning (CFL) (Chalupka et al., 2014, 2016), which achieves a CMR by merging values of a large observation space to yield discrete high-level variables taking values in a small finite set. Consider:
  • $\mathcal{L}$ has vector of endogenous variables $\mathbf{X}$ with range $\mathcal{X}$ and set of interventions $\mathcal{I}$,
  • $\mathcal{H}$ has vector of endogenous variables $\mathbf{Z}$ with range $\mathcal{Z}$ and set of interventions $\mathcal{J}$.
@@
A general framework for CMR is based on the notion of exact transformation (Rubenstein et al., 2017), which ensures interventional equivalence by matching the pushed forward low-level distributions to the high-level ones.
@@
A map $\tau:\mathcal{X}\to\mathcal{Z}$ is an exact transformation from $\mathcal{L}$ to $\mathcal{H}$ if it is surjective, and there exists a surjective map $\omega:\mathcal{I}\to\mathcal{J}$ such that for all $\mathbf{i}\in\mathcal{I}$
@@
The set of possible $\tau$ can be restricted to constructive transformations, which offer good interpretability and come with characterization results (Beckers and Halpern, 2019; Geiger et al., 2023a).
@@
$\tau:\mathcal{X}\to\mathcal{Z}$ is a constructive transformation between model $\mathcal{L}$ and $\mathcal{H}$ if there exists an alignment map $\pi$ relating indices of each high-level endogenous variable to a subset of indices of low-level endogenous variables such that for all $k\neq l$, $\pi(k)\cap\pi(n)=\emptyset$ and for each component $\tau_{k}$ of $\tau$ it exists a function $\bar{\tau}_{k}$ such that for all $\mathbf{x}$ in $\mathcal{X}$,
$\tau_{k}(\mathbf{x})=\bar{\tau}_{k}(\mathbf{x}_{\pi(k)})\,,$
i.e. each high-level variable $Z_{k}$ depends only on the subset of low-level variables given by $\pi(k)$.
The definition of a constructive transformation extends naturally to the intervention map $\omega$, for which we also require that interventions acting on high-level variable $k$ depend only on low-level interventions acting on variables in $\pi(k)$ (see App. A).
Reduction of a scientific model: illustration in mechanics.
We illustrate how CMR links to reduction of a scientific model using the above example of Fig. 0(a). Consider a system of points $\{\vec{OP}_{k}=(x_{k},y_{k})\}$ in a two-dimensional Galilean coordinate system $(O,\vec{i},\vec{j})$. If we are only interested in the position and speed of the system’s center of mass $G$ along the $x$-axis, the transformation to the high-level model is $\tau_{1}:(x_{1},y_{1},x_{2},...,v_{x,1},...))\mapsto\frac{1}{\sum_{k}m_{k}}\sum_{k}m_{k}x_{k}$ and $\tau_{2}:(x_{1},y_{1},x_{2},...,v_{x,1},...))\mapsto\frac{1}{\sum_{k}m_{k}}\sum_{k}m_{k}v_{x,k}$. Moreover, low-level interventions involving forces applied to each point mass solely affect the trajectory of the center of mass through the $x$ component of the net force, computed as the sum of all elementary forces, which corresponds to the mapping $\omega$ (see Fig. 0(a)). While there is no SCM associated to this system in the usual sense, there is one associated to the Euler numerical scheme simulating it (see App. A).
@@ 3 THEORETICAL ANALYSIS
As described in Fig. 0(b), we consider endogenous variables of a low-level model gathered in a (high-dimensional) random vector $\mathbf{X}$. A target scalar variable $Y=\tau_{0}(\mathbf{X})$ quantifies a property of interest of this model, and can be thought of as quantifying the presence or magnitude of a pattern or phenomenon in the data, using detector $\tau_{0}$. To generate a high-level causal explanation of this pattern, we learn a high-level SCM with fixed effect variable $Y$ and $n$ causes of $Y$ represented by a low-dimensional vector ${\mathbf{Z}}$, leading to a high-level (Markovian) causal graph $\mathcal{G}_{\mathcal{H}}$ with only arrows pointing from each $Z_{k}$ root node to leaf node $Y$. The low-level variables $\mathbf{X}$ are approximately mapped to the high-level one using a constructive transformation with alignement $\pi$. Approximate interventional consistency between models is obtained by defining a mapping $\omega$ between interventions at the low level, and their high-level counterpart.
@@ 3.1 TCR framework
(1) A low-level SCM $\mathcal{L}$ with $N$ endogenous variables $\{X_{1},X_{2},\dots,X_{N}\}$ and corresponding exogenous variables $\{U_{k}\}_{k=1..N}$ equipped with joint distribution $P(\mathbf{U})$. A set of low-level interventions parameterized by vector $\mathbf{i}\in\mathcal{I}$ with distribution $P(\mathbf{i})$, with each component $i_{k}$ affecting a unique endogenous variable $X_{k}$. Unintervened variables are assigned a fixed intervention parameter $i_{k}=0$. We only assume we can query (samples form) unintervened and interventional distributions of $\mathcal{L}$.
(2) A class of high-level SCMs $\{\mathcal{H}_{\bm{\gamma}}\}_{\gamma\in\Gamma}$ with (n+1) endogenous variables $\{Y,{Z}_{1},\dots,{Z}_{n}\}$ and associated exogenous variables $\{R_{k}\}_{k=0..n}$, equipped with a factorized distribution $P(\mathbf{R})=\prod P_{R_{k}}$. A set of high-level interventions parametrized by vector $\mathbf{j}\in\mathcal{J}$, with each component $j_{k}$ affecting a single node $Z_{k}$. Contrary to the (fixed) low-level model, the high-level model parameters $\bm{\gamma}$ need to be learned.
These two models are linked by a constructive transformation with two deterministic surjective maps $\tau$ and $\omega$ from low- to high-level endogenous variables and interventions, respectively, which decompose as

### v2

# Targeted Reduction of Causal Models

@@ Abstract
Why does a phenomenon occur? Addressing this question is central to most scientific inquiries and often relies on simulations of scientific models. As models become more intricate, deciphering the causes behind phenomena in high-dimensional spaces of interconnected variables becomes increasingly challenging. Causal Representation Learning (CRL) offers a promising avenue to uncover interpretable causal patterns within these simulations through an interventional lens. However, developing general CRL frameworks suitable for practical applications remains an open challenge. We introduce Targeted Causal Reduction (TCR), a method for condensing complex intervenable models into a concise set of causal factors that explain a specific target phenomenon. We propose an information theoretic objective to learn TCR from interventional data of simulations, establish identifiability for continuous variables under shift interventions and present a practical algorithm for learning TCRs. Its ability to generate interpretable high-level explanations from complex models is demonstrated on toy and mechanical systems, illustrating its potential to assist scientists in the study of complex phenomena in a broad range of disciplines.1111Code is available at: https://github.com/akekic/targeted-causal-reduction.git.
@@ 1 INTRODUCTION
Numerical models are indispensable in science for simulating real-world systems and generating etiological explanations—identifying the causes of specific phenomena. General circulation models, for example, shed light on the causes of global warming (Grassl, 2000), while computational brain models explore the origins of neurological pathologies (Breakspear, 2017; Deco and Kringelbach, 2014). These examples illustrate the increasing complexity of numerical scientific models, designed to faithfully capture the large number of mechanisms at play in these systems. However, this complexity comes at a cost: expanding parameter spaces and heightened computational demands. This trend, in turn, impacts the ability to generate high-level explanations, understandable by scientists and decision makers (Reichstein et al., 2019; Safavi et al., 2023).
Effective human explanations are often based on understanding a few causal relations between a limited number of variables. While the simulation of complex systems might rely on numerous simple mechanisms, extracting overarching causal relations between fewer relevant high-level variables remains largely unaddressed. In particular, while causal representation learning tries to explain data based on a learned latent causal graph (Wendong et al., 2023; Squires et al., 2023; von Kügelgen et al., 2023a), it currently has theoretical and practical limitations. CRL largely relies on preserving all information in the data to provide recoverability guaranties for the latent causes, while the idea of a high-level representation is precisely to discard irrelevant data.
In contrast, Causal Model Reduction (CMR), which aims to map a low-level causal model to a simpler high-level model with fewer or lower-dimensional variables, embraces the purpose of eliminating irrelevant information. However, existing CMR approaches, such as causal abstractions (Geiger et al., 2023; Zennaro et al., 2023) and Causal Feature Learning (CFL) (Chalupka et al., 2016) are not well-suited to causally describe many scientific models: they use discrete variables and typically rely on hard interventions, disconnecting causal variables from their parents. The following example shows, however, that simpler high-level causal models for continuous variables and soft interventions are natural and useful in domains such as physics.
Figure 1: Targeted Causal Reduction. (a) Example targeted model reduction: a model of the dynamics of a system of point masses connected by springs can be reduced to the trajectory of its center of mass. (b) Overview of TCR. Low-level variables $\mathbf{X}$ (simulation) are mapped to high-level variables $(\mathbf{Z},Y)$ with a fixed causal structure. The target $Y$ is known, while the causes $\mathbf{Z}$ and the high-level causal mechanism are learned. Additionally, we learn a mapping from low-level shift interventions $\mathbf{i}$ to high-level shift interventions ${\bm{\omega}}(\mathbf{i})$.
Consider a system of point masses connected by springs shown in Fig. 1(a), where each mass is influenced by random external forces. Its trajectory under the intervention of external forces can be accurately predicted by simulating the coupled equations of motion of individual point masses. However, if we are only interested in a particular macroscopic “target” variable of this system: the horizontal speed of the system’s center of mass at the end of an experiment, a key result from classical physics is that its motion will depend only on the sum of all horizontal components of external forces applied over time. We thus obtain a form of CMR: a much simpler system that accurately accounts for the effect of interventions in the system on the target variable.
This highlights the core elements needed for CMR in scientific models: (1) there is a clearly defined macroscopic target variable, (2) continuous low-level variables are reduced to a smaller set of continuous high-level variables, and (3) low-level interventions are soft: exerted forces modify the future trajectory but do not suppress the influence of other factors such as the past state of the system. Moreover, it is the combination of many low-level soft interventions across time that corresponds to a relevant high-level intervention. These three aspects are commonly found characteristics of studied real-world systems, motivating the development of CMR algorithms adapted to this setting.
In this paper, we introduce Targeted Causal Reduction (TCR), depicted in Fig. 1(b), a novel approach designed to simplify complex low-level models into high-level models, focused on explaining causal influences on an observable target variable $Y$. The key signal we use for learning are interventions applied to the low-level variables, which are mapped to high-level interventions in a way that captures the causal influences on $Y$ in a concise and interpretable way. We formulate this learning objective as a Kullback-Leibler divergence between the fitted high-level interventional model and the reduced low-level interventional distribution, leading to a practical learning algorithm for the case of linear reductions. Applications to high-dimensional synthetic and scientific models demonstrates accuracy and interpretability of our approach. We refer to the appendix for more related work (A) and proofs (C).
@@ 2.1 Structural causal models
Causal dependencies between variables can be described using Structural Causal Models (SCM) (Peters et al., 2017).
Notation.
We use boldface for column vectors, and $\mathbf{i}_{S}$ for the subvector of $\mathbf{i}$ restricted to the components in set $S$.
@@
  • a joint distribution $P_{\mathbf{U}}$ over exogenous variables $\{U_{j}\}_{j\leq n}$,
  • a directed graph $\mathcal{G}$ with $n$ vertices,
@@
such that for almost every $\mathbf{u}$, the system $\{x_{j}\coloneqq f_{j}(\textbf{pa}_{j},u_{j})\}$ has a unique solution $\mathbf{x}=\mathbf{g}(\mathbf{u})$, with $\mathbf{g}$ measurable.
The unique solvability condition is included in this definition because we consider a general class of SCMs by allowing cycles, that is, $\mathcal{G}$ may not be a DAG. Moreover, we allow hidden confounding through the potential lack of independence between the exogenous variables $\{U_{j}\}$. See Bongers et al. (2021) for a thorough study of these models. Under these conditions, the distribution $P_{\mathbf{U}}$ entails a well-defined joint distribution over the endogenous variables $P(\mathbf{X})$.
Interventions in SCMs involve replacing one or more structural equations, potentially modifying exogenous distributions, and adding or removing arrows in the original graph to reflect changes in dependencies between variables. An intervention transforms the original model $\mathcal{M}=(\mathcal{G},\mathbb{S},P_{\mathbf{U}})$ into an intervened model $\mathcal{M}^{(\mathbf{i})}=(\mathcal{G}^{(\mathbf{i})},\mathbb{S}^{(\mathbf{i})},P_{\mathbf{U}}^{(\mathbf{i})})$, where $\mathbf{i}$ is the vector parameterizing the intervention. The base probability distribution of the unintervened model is denoted $P_{\mathcal{M}}^{(0)}(\mathbf{X})$ or simply $P_{\mathcal{M}}(\mathbf{X})$ and the interventional distribution associated with $\mathcal{M}^{(\mathbf{i})}$ is denoted $P_{\mathcal{M}}^{(\mathbf{i})}(\mathbf{X})$.
Classical $do$-interventions set a structural equation to a constant, removing all influences of the parents on the intervened variable. This can be problematic for studying how the influence of low-level variables is propagated to the target, since for simultaneous interventions, the effects of some interventions can be masked by others. The probability of such masking increases as the number of low-level variables grows. Soft interventions, on the other hand, modify an equation while keeping the set of parents unchanged. This is more appropriate in our setting, since it propagates the information from all interventions to the target simultaneously.
Large classes of soft interventions can be designed to match domain knowledge (Besserve and Schölkopf, 2022). Notably, shift interventions modify the structural equation of endogenous variable $l$ through shifting it by a scalar parameter $i$
@@
2.2 Simulations and Causal Models
We use the term scientific model to refer to a generative model that relies on a set of equations to represent a phenomenon. What distinguishes such models from generative models in machine learning is their decomposability into elementary functions, encoding domain knowledge about the mechanisms being investigated. Simulations based on the numerical solution of scientific models can often be expressed as SCMs; This notably includes Ordinary (ODE) (Mooij et al., 2013) and Stochastic Differential Equations (SDE) (Hansen and Sokol, 2014). A simulator can thus be seen as a low-level causal model, from which samples of unintervened and intervened distributions can be generated. This forms the basis of the causal framework for learning high-level explanations for simulators developed in Sec. 3.1.
@@ 2.3 Causal Model Reductions (CMR)
We consider as CMR any (possibly approximate) mapping from a low-level SCM $\mathcal{L}$ to a simpler high-level SCM $\mathcal{H}$. An example is CFL (Chalupka et al., 2015, 2016), which achieves a CMR by merging values of a large observation space to yield discrete high-level variables taking values in a small finite set. Consider:
  • $\mathcal{L}$ has a vector of endogenous variables $\mathbf{X}$ with range $\mathcal{X}$ and a set of interventions $\mathcal{I}$,
  • $\mathcal{H}$ has a vector of endogenous variables $\mathbf{Z}$ with range $\mathcal{Z}$ and a set of interventions $\mathcal{J}$.
@@
A general framework for CMR is based on the notion of exact transformation, which ensures interventional consistency by matching the push-forward low-level distributions to the high-level ones.
@@
A map $\tau:\mathcal{X}\to\mathcal{Z}$ is an exact transformation from $\mathcal{L}$ to $\mathcal{H}$ if it is surjective, and there exists a surjective intervention map $\omega:\mathcal{I}\to\mathcal{J}$ such that for all $\mathbf{i}\in\mathcal{I}$
@@
The set of possible $\tau$ can be restricted to constructive transformations, where high-level variables depend only on non-overlapping subsets of low-level variables. This eases interpretability of CMR and comes with characterization results (Beckers and Halpern, 2019; Geiger et al., 2023).
@@
$\tau:\mathcal{X}\to\mathcal{Z}$ is a constructive transformation between model $\mathcal{L}$ and $\mathcal{H}$ if there exists an alignment map $\pi$ relating indices of each high-level endogenous variable to a subset of indices of low-level endogenous variables such that for all $k\neq l$, $\pi(k)\cap\pi(l)=\emptyset$ and for each component $\tau_{k}$ of $\tau$ it exists a function $\bar{\tau}_{k}$ such that for all $\mathbf{x}$ in $\mathcal{X}$,
$\tau_{k}(\mathbf{x})=\bar{\tau}_{k}(\mathbf{x}_{\pi(k)})\,.$
The intervention map $\omega$ of constructive exact transformations are required to be constructive as well, such that acting on high-level variable $k$ depends only on low-level interventions acting on variables in $\pi(k)$ (see App. B).
@@ 3 THEORETICAL ANALYSIS
As described in Fig. 1(b), we consider endogenous variables of a low-level model gathered in a (high-dimensional) random vector $\mathbf{X}$. A target scalar variable $Y=\tau_{0}(\mathbf{X})$ quantifies a property of interest of this model, and can be thought of as quantifying the presence or magnitude of a phenomenon in the data, using detector $\tau_{0}$. To generate a high-level causal explanation of this phenomenon, we learn a high-level SCM with a fixed causal structure, where the known effect variable $Y$ is caused by $n$ learned independent high-level variables $Z_{k}$. The low-level variables $\mathbf{X}$ are approximately mapped to the high-level variables ${\mathbf{Z}}$ using a constructive transformation and an associated constructive interventional map with the same alignment $\pi$.
@@ 3.1 TCR framework
(1) A low-level SCM $\mathcal{L}$ with $N$ endogenous variables $\{X_{1},\dots,X_{N}\}$ and corresponding exogenous variables $\{U_{k}\}_{k=1..N}$ equipped with joint distribution $P(\mathbf{U})$. A set of low-level shift interventions parameterized by vector $\mathbf{i}{\in}\mathcal{I}$ with distribution $P(\mathbf{i})$, with each component $i_{k}$ acting on one endogenous variable $X_{k}$. We only assume we can sample from unintervened and interventional distributions of $\mathcal{L}$.
(2) A class of high-level SCMs $\{\mathcal{H}_{\bm{\gamma}}\}_{\gamma\in\Gamma}$ with (n+1) endogenous variables $\{Y,{Z}_{1},\dots,{Z}_{n}\}$ and associated exogenous variables $\{R_{k}\}_{k=0..n}$, equipped with a factorized distribution $P(\mathbf{R})=\prod P_{R_{k}}$. A set of high-level shift interventions parametrized by vector $\mathbf{j}{\in}\mathcal{J}$, with each component $j_{k}$ affecting a single node $Z_{k}$. In contrast to the (fixed) low-level model, the high-level model parameters $\bm{\gamma}$ are learned.
These two levels are linked by a constructive transformation with two deterministic surjective maps $\tau$ and $\omega$ from low- to high-level endogenous variables and interventions, respectively, which decompose as
@@
A high-level model involves the following mechanisms, which need to be learned: (1) The marginal distribution of each high-level cause $P^{(\mathbf{j})}({Z}_{k})$ in all high-level interventional settings $\mathbf{j}$. (2) The mechanism $P(Y|\mathbf{Z})$ mapping high-level causes to $Y$, comprised of the distribution of the exogenous variable $R_{0}$ and the structural equation